OpenStudy (anonymous):

In 1990, the population of the Midwest was about 60 million. During the 1990s, the population of this area increased an average of about 0.4 million per year. The population of the West was about 53 million in 1990. The population of this area increased an average of about 1 million per year during the 1990s. Assume that the rate of growth of these areas remains the same. Estimate when the population of the West would be equal to the population of the Midwest.

OpenStudy (math&ing001):

Let x be the number of years that's gonna take for the two populations to be equal. 60+0.4x=53+x Solve for x. The year when the two populations to be equal will be 1990+x.

OpenStudy (anonymous):

I do not get it

OpenStudy (math&ing001):

Did you study how to solve a linear equation ?

OpenStudy (anonymous):

X = 11.6667

OpenStudy (math&ing001):

Yep ! So the population of the West would be equal to the population of the Midwest at around 2002.

OpenStudy (anonymous):

do you write it like this 60 + 0.4 x = 53 + x Solve for x. 1990+x X = 11.6667

OpenStudy (math&ing001):

Yeah, all I did after was 1990+x=1990+11.6667=2001.6667 And because it's a year we round it up to 2002.